# Comment on “Neutrino interaction with matter in a noninertial frame”

### v1

@@ IV Final remarks
In this comment, we obtain the complete energy levels for Dvornikov’s paper [2], that is, the energy levels dependent on two quantum numbers, namely, the radial quantum number (given by $N=0,1,2,\ldots$) and the angular quantum number (given by $J_{z}=\pm\frac{1}{2},\pm\frac{3}{2},\pm\frac{5}{2},\ldots$). In particular, what motivated us to do this was the fact that the quantized energy levels for particles (fermions or bosons) in polar, cylindrical, or spherical coordinates (i.e. 2D and 3D systems) depend on two quantum numbers: a radial quantum number and an angular quantum number. In view of this, the following question/doubt arose: Why do the energy levels in Dvornikov’s paper only depend on one quantum number? In other words, Where did the angular quantum number given by $J_{z}$ go? So, using Studenikin’s paper as a starting point (as well as others in the literature [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 15, 16]), we write one of the equations from Dvornikov’s paper [2] in a matrix form. Later, we use the four-component Dirac spinor and obtain a set/system of four coupled first-order differential equations. From the first two equations with $m\simeq 0$, we obtain a (compact) second-order differential equation for two spinor components (i.e. for $\psi_{3}$ and $\psi_{4}$). So, solving this equation, we obtain the neutrino energy levels, which explicitly depend on both $N$ and $J_{z}$. Finally, we note that for $J_{z}>0$ (positive angular momentum) with $u=+1$ (component $\psi_{3}$), we obtain exactly the particular energy levels of Dvornikov’s paper [2].

### v2

@@ IV Final remarks
In this comment, we obtain the complete energy levels for Dvornikov’s paper [2], that is, the energy levels dependent on two quantum numbers, namely, the radial quantum number (given by $N=0,1,2,\ldots$) and the angular quantum number (given by $J_{z}=\pm\frac{1}{2},\pm\frac{3}{2},\pm\frac{5}{2},\ldots$). In particular, what motivated us to do this was the fact that the quantized energy levels for particles (fermions or bosons) in polar, cylindrical, or spherical coordinates depend on two quantum numbers: a radial quantum number and an angular quantum number. In view of this, the following question/doubt arose: Why do the energy levels in Dvornikov’s paper only depend on one quantum number? In other words, Where did the angular quantum number given by $J_{z}$ go? So, using Studenikin’s paper as a starting point (as well as others in the literature [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 15, 16]), we write one of the equations from Dvornikov’s paper [2] in a matrix form. Later, we use the four-component Dirac spinor and obtain a set/system of four coupled first-order differential equations. From the first two equations with $m\to 0$, we obtain a (compact) second-order differential equation for the last two spinor components (i.e. $\psi_{3}$ and $\psi_{4}$). So, solving this equation, we obtain the neutrino energy levels, which explicitly depend on both $N$ and $J_{z}$. Finally, we note that for $J_{z}>0$ (positive angular momentum) with $u=+1$ (component $\psi_{3}$), we obtain exactly the particular energy levels of Dvornikov’s paper [2].